OpenStudy (anonymous):

In a group of 100 persons, 85 had Rh positive blood (Rh factor present), 45 had Type O blood, and 38 had both type O and Rh positive blood. If one person had been selected at random from the group, what would the probability be: that his blood type was Type O and Rh+? that his blood type was not O? that his blood type was Rh-? that his blood type was Type O and Rh-? that his blood type was Rh+? that his blood type was not Type O or else was Rh+?

OpenStudy (amistre64):

something to do with: P(A) + P(B) - P(AnB) seems familiar

OpenStudy (amistre64):

P(AuB) = .... when A and B are not independant

OpenStudy (amistre64):

or the more brute method is to venn it up

OpenStudy (amistre64):



OpenStudy (amistre64):

since it asks for probability, you would need to convert those %s to decimals

OpenStudy (anonymous):

Sorry, I was away from the computer. & I'm not really sure how to do that? I just don't understand how it adds up to 100 because its 85, 45, and 38 and that isn't even

OpenStudy (amistre64):

notice that the intersection, people with BOTH O and Rh is 38 this contributes to both groups of O and Rh x + 38 = 85 y + 38 = 45

OpenStudy (amistre64):

x = 47, y = 7 are the areas of the corresponding venn diagram

OpenStudy (amistre64):

since 7+47+38 = 92, that leaves 8 left over that are neither

OpenStudy (amistre64):

if we simply make the mistake and add up 85 and 45 x + 38 = 85 y + 38 = 45 x + 38 + y + 38 ... we have counted the "38" twice, so we need to subtract out a 38 85 + 45 - 38 is the total of the sets ....

OpenStudy (anonymous):

Oh, okay. I see that now. So when I'm trying to find the fractions do I use the numbers they gave me or the ones that you have found out?

OpenStudy (amistre64):

id use the numbers i found out, do you have to make fractions? the decimals are simple enough

OpenStudy (anonymous):

Yes I have to use fractions.. So for like that his blood type was Type O and Rh+? what would I do?

OpenStudy (amistre64):



OpenStudy (amistre64):

O and Rh was already given in the setup .... it is also located at the intersection of the 2 circles: .38

OpenStudy (amistre64):

notO is all the values that are NOT in O: .47 + .08

OpenStudy (amistre64):

that his blood type was Rh-? are all the values that are NOT in Rh: .07+.08

OpenStudy (anonymous):

Oh okay, so it's not super complicated, its really just adding the proper numbers?

OpenStudy (amistre64):

correct blood type was Type O and Rh- ? everyting in O that is NOT in Rh: .07

OpenStudy (anonymous):

Oh okay, I see now. Well thank you so much for your help :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

I just checked my answers & I got them all right! I started out using decimals and then converted them to fractions. It was way easier that way. Thanks again!